# Statistical Learning of Distributionally Robust Stochastic Control in Continuous State Spaces

### v1

@@ Appendix C Proofs for Section 3
By Proposition 2.1, it is easy to see that $\mathbf{H}$ is also a $\alpha$-contraction on $U_{b}(\mathbb{X})$, with 00 as its unique fixed point.
@@ C.2 Proof of Theorem 1
We remark that our proof techniques have similarities with that in Lee and Raginsky [2018]. By Proposition 3.1, to achieve an upper bound on the uniform learning error, it suffices to prove an upper bound for $\left\|\mathbf{T}(u^{*})-\mathcal{T}(u^{*})\right\|$.
To do this, we first rewrite the Bellman operator using its dual form. By strong duality [Blanchet and Murthy, 2019], for $\mathcal{P}=\left\{{\mu\in\mathcal{P}(\mathcal{X}):W(\mu,\mu_{0})\leq\delta}\right\}$
$\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}u^{*}(f(z,w))\psi(dw)=\sup_{\lambda\geq 0}-\lambda\delta+\int_{\mathbb{X}}\inf_{y\in\mathbb{W}}\left[u^{*}(f(z,y))+\lambda c(w,y)\right]\mu_{0}(dw).$
@@
$\int_{\mathbb{X}}\inf_{y\in\mathbb{W}}\left[u^{*}(f(z,y))+\lambda c(w,y)\right]\mu_{0}(dw)\leq\int_{\mathbb{X}}u^{*}(f(z,w))\mu_{0}(dw)\leq\left\|u\right\|$
@@
$\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}u^{*}(f(z,w))\psi(dw)\geq 0,$
it suffices to maximize $\lambda$ over $\Lambda:=[0,\delta^{-1}\left\|u^{*}\right\|]$.
@@
$\leq\alpha\sup_{z\in\mathbb{X}\times\mathbb{A}}\left|\sup_{\lambda\in\Lambda}\left[\int_{\mathbb{X}}\inf_{y\in\mathbb{W}}\left[u^{*}(f(z,y))+\lambda c(w,y)\right]\mu_{0}(dw)-\lambda\delta\right]-\sup_{\lambda\in\Lambda}\left[\int_{\mathbb{X}}\inf_{y\in\mathbb{W}}\left[u^{*}(f(z,y))+\lambda c(w,y)\right]\hat{\mu}(dw)-\lambda\delta\right]\right|$
$\leq\alpha\sup_{z\in\mathbb{X}\times\mathbb{A}}\sup_{\lambda\in\Lambda}\left|\int_{\mathbb{X}}\inf_{y\in\mathbb{W}}\left[u^{*}(f(z,y))+\lambda c(w,y)\right](\mu_{0}-\hat{\mu})(dw)\right|$
$=\alpha\sup_{\theta\in\Theta}\left|(\mu_{0}-\hat{\mu})[g_{\theta}]\right|$
@@
$g_{\theta}=\inf_{y\in\mathbb{W}}\left[u^{*}(f(z,y))+\lambda c(\cdot,y)\right].$
@@
$R_{n}(\mathbf{w},g_{\theta}):=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\epsilon_{i}g_{\theta}(w_{i})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\epsilon_{i}\inf_{w\in\mathbb{W}}\left[u^{*}(f(z,y))+\lambda c(w_{i},y)\right].$ (C.1)
@@
From empirical process theory, see e.g. Wainwright [2019, Theorem 4.10], w.p. at least $1-\eta$,
@@
where $g_{\vee}:=\sup_{g\in\mathcal{G}}\|g\|\leq\left\|u^{*}\right\|\leq\beta$. Thus, we proceed to bound $\mathcal{R}_{n}(\mathbf{w},\mathcal{G})$ and hence $\mathcal{R}_{n}(\mathcal{G})$. We achieve this by using subgaussian processes and entropy integrals.
@@
$=E_{\epsilon}\exp\left(\frac{\xi}{\sqrt{n}}\sum_{i=1}^{n}\epsilon_{i}\left[\inf_{y\in\mathbb{W}}\left[u^{*}(f(z,y))+\lambda c(w_{i},y)\right]-\inf_{y\in\mathbb{W}}\left[u^{*}(f(z^{\prime},y))+\lambda^{\prime}c(w_{i},y)\right]\right]\right)$
$\leq\exp\left(\frac{\xi^{2}}{2n}\sum_{i=1}^{n}\left[\inf_{y\in\mathbb{W}}\left[u^{*}(f(z,y))+\lambda c(w_{i},y)\right]-\inf_{y\in\mathbb{W}}\left[u^{*}(f(z^{\prime},y))+\lambda^{\prime}c(w_{i},y)\right]\right]^{2}\right)$
$\leq\exp\left(\frac{\xi^{2}}{2n}\sum_{i=1}^{n}\sup_{y\in\mathbb{W}}\left|u^{*}(f(z,y))-u^{*}(f(z^{\prime},y))+(\lambda^{\prime}-\lambda)c(w_{i},y)\right|^{2}\right)$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\exp\left(\frac{\xi^{2}}{2}\left(\sup_{y\in\mathbb{W}}\left|u^{*}(f(z,y))-u^{*}(f(z^{\prime},y))\right|+\left|\lambda-\lambda^{\prime}\right|c_{\vee}\right)^{2}\right)$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\exp\left(\frac{\xi^{2}}{2}\left(\ell(\left|x-x^{\prime}\right|+|a-a^{\prime}|)+\left|\lambda-\lambda^{\prime}\right|c_{\vee}\right)^{2}\right)$
@@
$\rho(\theta,\theta^{\prime}):=\ell(|x-x^{\prime}|+|a-a^{\prime}|)+c_{\vee}|\lambda-\lambda^{\prime}|,$
@@
Therefore, using Dudley’s entropy integral for subgaussian processes [Wainwright, 2019, Chapter 5], the empirical Rademacher complexity in (C.2) can be bounded by
@@
[…]
@@
$D_{\vee}$ $:=\ell(\mathrm{diam}(\mathbb{X})+\mathrm{diam}(\mathbb{A}))+c_{\vee}\delta^{-1}\beta+1$
$\geq\ell(\mathrm{diam}(\mathbb{X})+\mathrm{diam}(\mathbb{A}))+c_{\vee}\mathrm{diam}(\Lambda)+1$
@@
[…]
@@
$\mathcal{N}(\epsilon;\Theta,\rho)$ $=\mathcal{N}(\epsilon;\mathbb{X}\times\mathbb{A},\ell|\cdot|)\cdot\mathcal{N}(\epsilon;\Lambda,c_{\vee}|\cdot|)$
$=\mathcal{N}(\epsilon;\mathbb{X},\ell|\cdot|)\cdot\mathcal{N}(\epsilon;\mathbb{A},\ell|\cdot|)\cdot\mathcal{N}(\epsilon;\Lambda,c_{\vee}|\cdot|)$
$\leq\left(1+\frac{\ell\mathrm{diam}(\mathbb{X})}{\epsilon}\right)^{d_{\mathbb{X}}}\left(1+\frac{\ell\mathrm{diam}(\mathbb{A})}{\epsilon}\right)^{d_{\mathbb{A}}}\left(1+\frac{c_{\vee}\mathrm{diam}(\Lambda)}{\epsilon}\right).$
@@
[…]
@@
$\leq\alpha\beta\sup_{\theta\in\Theta}\left|(\mu_{n}-\mu_{0})[g_{\theta}]\right|$
$\stackrel{{\scriptstyle(i)}}{{\leq}}2\alpha\beta\mathcal{R}_{n}(\mathcal{G})+\alpha\beta\sqrt{\frac{2g_{\vee}}{n}\log\left(\frac{1}{\eta}\right)}$
$\leq\left(32\sqrt{\pi}\alpha\beta D_{\vee}\sqrt{d_{\mathbb{X}}+d_{\mathbb{A}}+1}+\sqrt{2}\alpha\beta^{3/2}\sqrt{\log\left(\frac{1}{\eta}\right)}\right)n^{-\frac{1}{2}}$
where $(i)$ holds w.p. at least $1-\eta$.
@@ C.3 Proof of Theorem 2
$\left\|\mathbf{T}(u^{*})-\mathcal{T}(u^{*})\right\|\leq\alpha\sup_{z\in\mathbb{X}\times\mathbb{A}}\left|\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}u^{*}(f(z,w))\psi(dw)-\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}u^{*}(f(z,w))\psi(dw)\right|$
Define the function class :=$\mathcal{F}:=\left\{{u^{*}(f(z,\cdot)):z\in\mathbb{X}\times\mathbb{A})}\right\}$. By Duchi and Namkoong [2021, Corollary 1], the r.h.s. satisfies w.p. at least $1-2\mathcal{N}(\epsilon/3;\mathcal{F},\left\|\cdot\right\|)e^{-t}$,
$\sup_{z\in\mathbb{X}\times\mathbb{A}}\left|\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}u^{*}(f(x,a,w))\psi(dw)-\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}u^{*}(f(x,a,w))\psi(dw)\right|\leq 30\beta\epsilon$
@@
$\epsilon=n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\rho)^{2}\left(\frac{c_{k}(\rho)}{c_{k}(\rho)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{t+2\log n}\right).$
@@
$\left\|\mathbf{T}(u^{*})-\mathcal{T}(u^{*})\right\|\leq 30\alpha\beta n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\rho)^{2}\left(\frac{c_{k}(\rho)}{c_{k}(\rho)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{\log(2\mathcal{N}(\epsilon/3;\mathcal{F},\left\|\cdot\right\|)+\log\frac{1}{\eta}+2\log n}\right).$
By the uniform Lipschitz assumption of $z\rightarrow u^{*}(f(z,\cdot))$ and that $\epsilon\geq n^{-1/2}$ for all $t\geq 1$, we have that by van der Vaart and Wellner [1996, Chapter 2.7.4]
$\log\mathcal{N}(\epsilon/3;\mathcal{F},\left\|\cdot\right\|)$ $\leq\log\mathcal{N}\left(\frac{\epsilon\wedge 1}{3\ell};\mathbb{X}\times\mathbb{A},\left|\cdot\right|\right)$
$\leq d_{\mathbb{X}}\log\left(1+\frac{3\ell\mathrm{diam}(\mathbb{X})}{\epsilon\wedge 1}\right)+{d_{\mathbb{A}}}\log\left(1+\frac{3\ell\mathrm{diam}(\mathbb{A})}{\epsilon\wedge 1}\right)$
$\leq d_{\mathbb{X}}\log\left(1+3\ell\mathrm{diam}(\mathbb{X})\right)+{d_{\mathbb{A}}}\log\left(1+3\ell\mathrm{diam}(\mathbb{A})\right)+\frac{1}{2}(d_{\mathbb{X}}+d_{\mathbb{A}})\log n.$
Therefore, defining $D=d_{\mathbb{X}}\log\left(1+3\ell\mathrm{diam}(\mathbb{X})\right)+{d_{\mathbb{A}}}\log\left(1+3\ell\mathrm{diam}(\mathbb{A})\right)$, we conclude that by Proposition 3.1, the estimation error
@@
$\stackrel{{\scriptstyle(i)}}{{\leq}}30\beta^{2}n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\rho)^{2}\left(\frac{c_{k}(\rho)}{c_{k}(\rho)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{D+\log\frac{1}{\eta}+2(d_{\mathbb{X}}+d_{\mathbb{A}})\log n}\right)$

### v2

@@ Appendix C Proofs for Section 4
By Proposition A.1, it is easy to see that $\mathbf{H}$ is also a $\alpha$-contraction on $U_{b}(\mathbb{X})$, with 0 as its unique fixed point.
@@ C.2 Proof of Theorem 2
We remark that our proof techniques have similarities with that in Lee and Raginsky, [22]. By Proposition 4.1, to achieve an upper bound on the uniform learning error, it suffices to prove an upper bound for $\left\|\mathbf{T}(u^{*})-\mathcal{T}(u^{*})\right\|$.
To do this, we first rewrite the Bellman operator using its dual form. By strong duality [5], for $\mathcal{P}=\left\{{\mu\in\mathcal{P}(\mathcal{X}):W(\mu,\mu_{0})\leq\delta}\right\}$
$\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}[r(z,w)+\alpha u^{*}(f(z,w))]\psi(dw)=\sup_{\lambda\geq 0}-\lambda\delta+\int_{\mathbb{W}}\inf_{y\in\mathbb{W}}\left[r(z,y)+\alpha u^{*}(f(z,y))+\lambda c(w,y)\right]\mu_{0}(dw).$
@@
$\int_{\mathbb{W}}\inf_{y\in\mathbb{W}}\left[r(z,y)+\alpha u^{*}(f(z,y))+\lambda c(w,y)\right]\mu_{0}(dw)\leq\int_{\mathbb{W}}[r(z,w)+\alpha u^{*}(f(z,w))]\mu_{0}(dw)\leq 1+\alpha\left\|u^{*}\right\|$
@@
$\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}r(z,w)+\alpha u^{*}(f(z,w))\psi(dw)\geq 0,$
it suffices to maximize $\lambda$ over $\Lambda:=[0,\delta^{-1}(1+\alpha\left\|u^{*}\right\|)]$.
@@
$\leq\sup_{z\in\mathbb{X}\times\mathbb{A}}\left|\sup_{\lambda\in\Lambda}\left[\int_{\mathbb{W}}\inf_{y\in\mathbb{W}}\left[r(z,y)+\alpha u^{*}(f(z,y))+\lambda c(w,y)\right]\mu_{0}(dw)-\lambda\delta\right]-\sup_{\lambda\in\Lambda}\left[\int_{\mathbb{W}}\inf_{y\in\mathbb{W}}\left[r(z,y)+\alpha u^{*}(f(z,y))+\lambda c(w,y)\right]\hat{\mu}(dw)-\lambda\delta\right]\right|$
$\leq\sup_{z\in\mathbb{X}\times\mathbb{A}}\sup_{\lambda\in\Lambda}\left|\int_{\mathbb{X}}\inf_{y\in\mathbb{W}}\left[r(z,y)+\alpha u^{*}(f(z,y))+\lambda c(w,y)\right](\mu_{0}-\hat{\mu})(dw)\right|$
$=\sup_{\theta\in\Theta}\left|(\mu_{0}-\hat{\mu})[g_{\theta}]\right|$
@@
$g_{\theta}=\inf_{y\in\mathbb{W}}\left[r(z,y)+\alpha u^{*}(f(z,y))+\lambda c(\cdot,y)\right].$
@@
$R_{n}(\mathbf{w},g_{\theta}):=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\epsilon_{i}g_{\theta}(w_{i})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\epsilon_{i}\inf_{y\in\mathbb{W}}\left[r(z,y)+\alpha u^{*}(f(z,y))+\lambda c(w_{i},y)\right].$ (C.1)
@@
From empirical process theory, see e.g. Wainwright, [45, Theorem 4.10], w.p. at least $1-\eta$,
@@
where $g_{\vee}:=\sup_{g\in\mathcal{G}}\|g\|\leq 1+\alpha\left\|u^{*}\right\|\leq 1+\alpha\beta=\beta$. Thus, we proceed to bound $\mathcal{R}_{n}(\mathbf{w},\mathcal{G})$ and hence $\mathcal{R}_{n}(\mathcal{G})$. We achieve this by using subgaussian processes and entropy integrals.
@@
$=E_{\epsilon}\exp\left(\frac{\xi}{\sqrt{n}}\sum_{i=1}^{n}\epsilon_{i}\left[\inf_{y\in\mathbb{W}}\left[r(z,y)+\alpha u^{*}(f(z,y))+\lambda c(w_{i},y)\right]-\inf_{y\in\mathbb{W}}\left[r(z^{\prime},y)+\alpha u^{*}(f(z^{\prime},y))+\lambda^{\prime}c(w_{i},y)\right]\right]\right)$
$\leq\exp\left(\frac{\xi^{2}}{2n}\sum_{i=1}^{n}\left[\inf_{y\in\mathbb{W}}\left[r(z,y)+\alpha u^{*}(f(z,y))+\lambda c(w_{i},y)\right]-\inf_{y\in\mathbb{W}}\left[r(z^{\prime},y)+\alpha u^{*}(f(z^{\prime},y))+\lambda^{\prime}c(w_{i},y)\right]\right]^{2}\right)$
$\leq\exp\left(\frac{\xi^{2}}{2n}\sum_{i=1}^{n}\sup_{y\in\mathbb{W}}\left|r(z,y)+\alpha u^{*}(f(z,y))-r(z^{\prime},y)-\alpha u^{*}(f(z^{\prime},y))+(\lambda^{\prime}-\lambda)c(w_{i},y)\right|^{2}\right)$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\exp\left(\frac{\xi^{2}}{2}\left(\sup_{y\in\mathbb{W}}\left|r(z,y)+\alpha u^{*}(f(z,y))-r(z^{\prime},y)-\alpha u^{*}(f(z^{\prime},y))\right|+\left|\lambda-\lambda^{\prime}\right|c_{\vee}\right)^{2}\right)$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\exp\left(\frac{\xi^{2}}{2}\left(L(\left|x-x^{\prime}\right|+|a-a^{\prime}|)+\left|\lambda-\lambda^{\prime}\right|c_{\vee}\right)^{2}\right)$
@@
$\rho(\theta,\theta^{\prime}):=L(|x-x^{\prime}|+|a-a^{\prime}|)+c_{\vee}|\lambda-\lambda^{\prime}|,$
@@
Therefore, using Dudley’s entropy integral for subgaussian processes [45, Chapter 5], the empirical Rademacher complexity in (C.2) can be bounded by
@@
$\leq\frac{32}{\sqrt{n}}\int_{0}^{D_{\vee}}\sqrt{\log\mathcal{N}(\epsilon;\Theta,\rho)}d\epsilon$
@@
$D_{\vee}$ $=L(\mathrm{diam}(\mathbb{X})+\mathrm{diam}(\mathbb{A}))+c_{\vee}\delta^{-1}\beta+1$
$\geq L(\mathrm{diam}(\mathbb{X})+\mathrm{diam}(\mathbb{A}))+c_{\vee}\mathrm{diam}(\Lambda)+1$
@@
$\mathcal{R}_{n}(\mathcal{G})=E_{\mu^{n}_{0}}\mathcal{R}_{n}(\mathbf{W},\mathcal{G})\leq\frac{32}{\sqrt{n}}\int_{0}^{D_{\vee}}\sqrt{\log\mathcal{N}(\epsilon;\Theta,\rho)}d\epsilon$
@@
$\mathcal{N}(\epsilon;\Theta,\rho)$ $=\mathcal{N}(\epsilon;\mathbb{X}\times\mathbb{A},L|\cdot|)\cdot\mathcal{N}(\epsilon;\Lambda,c_{\vee}|\cdot|)$
$=\mathcal{N}(\epsilon;\mathbb{X},L|\cdot|)\cdot\mathcal{N}(\epsilon;\mathbb{A},L|\cdot|)\cdot\mathcal{N}(\epsilon;\Lambda,c_{\vee}|\cdot|)$
$\leq\left(1+\frac{L\mathrm{diam}(\mathbb{X})}{\epsilon}\right)^{d_{\mathbb{X}}}\left(1+\frac{L\mathrm{diam}(\mathbb{A})}{\epsilon}\right)^{d_{\mathbb{A}}}\left(1+\frac{c_{\vee}\mathrm{diam}(\Lambda)}{\epsilon}\right).$
@@
$\int_{0}^{D_{\vee}}\sqrt{\log\mathcal{N}(\epsilon;\Theta,\rho)}d\epsilon$ $\leq\int_{0}^{D_{\vee}}\sqrt{(d_{\mathbb{X}}+d_{\mathbb{A}}+1)\log\left(1+\max\left\{{\frac{L)\mathrm{diam}(\mathbb{A})}{\epsilon},\frac{L\mathrm{diam}(\mathbb{X})}{\epsilon},\frac{c_{\vee}\mathrm{diam}(\Lambda)}{\epsilon}}\right\}\right)}d\epsilon$
$\leq\int_{0}^{D_{\vee}}\sqrt{(d_{\mathbb{X}}+d_{\mathbb{A}}+1)\log(D_{\vee}/\epsilon)}d\epsilon$
@@
$\leq\beta\sup_{\theta\in\Theta}\left|(\mu_{n}-\mu_{0})[g_{\theta}]\right|$
$\stackrel{{\scriptstyle(i)}}{{\leq}}2\beta\mathcal{R}_{n}(\mathcal{G})+\beta\sqrt{\frac{2g_{\vee}}{n}\log\left(\frac{1}{\eta}\right)}$
$\leq\left(32\sqrt{\pi}\beta D_{\vee}\sqrt{d_{\mathbb{X}}+d_{\mathbb{A}}+1}+\sqrt{2}\beta^{3/2}\sqrt{\log\left(\frac{1}{\eta}\right)}\right)n^{-\frac{1}{2}}$
where $(i)$ holds w.p. at least $1-\eta$. This bound implies the theorem as we note that $d_{\mathbb{X}}\geq 1$.
@@ C.3 Proof of Theorem 3
$\left\|\mathbf{T}(u^{*})-\mathcal{T}(u^{*})\right\|\leq\sup_{z\in\mathbb{X}\times\mathbb{A}}\left|\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}[r(z,w)+\alpha u^{*}(f(z,w))]\psi(dw)-\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}[r(z,w)+\alpha u^{*}(f(z,w))]\psi(dw)\right|$
Define the function class $\mathcal{F}:=\left\{{r(z,\cdot)+\alpha u^{*}(f(z,\cdot)):z\in\mathbb{X}\times\mathbb{A})}\right\}$. By Duchi and Namkoong, [9, Corollary 1], the r.h.s. satisfies w.p. at least $1-2\mathcal{N}(\epsilon/3;\mathcal{F},\left\|\cdot\right\|)e^{-t}$,
$\sup_{z\in\mathbb{X}\times\mathbb{A}}\left|\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}[r(z,w)+\alpha u^{*}(f(z,w))]\psi(dw)-\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}[r(z,w)+\alpha u^{*}(f(z,w))]\psi(dw)\right|\leq 30(1+\alpha\beta)\epsilon$
@@
$\epsilon=n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\delta)^{2}\left(\frac{c_{k}(\delta)}{c_{k}(\delta)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{t+2\log n}\right).$
@@
$\left\|\mathbf{T}(u^{*})-\mathcal{T}(u^{*})\right\|\leq 30(1+\alpha\beta)n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\delta)^{2}\left(\frac{c_{k}(\delta)}{c_{k}(\delta)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{\log(2\mathcal{N}(\epsilon/3;\mathcal{F},\left\|\cdot\right\|)+\log\frac{1}{\eta}+2\log n}\right).$
By the uniform Lipschitz assumptions and that $\epsilon\geq n^{-1/2}$ for all $t\geq 1$, we have that by van der Vaart and Wellner, [43, Chapter 2.7.4]
$\log\mathcal{N}(\epsilon/3;\mathcal{F},\left\|\cdot\right\|)$ $\leq\log\mathcal{N}\left(\frac{\epsilon\wedge 1}{3L};\mathbb{X}\times\mathbb{A},\left|\cdot\right|\right)$
$\leq d_{\mathbb{X}}\log\left(1+\frac{3L\mathrm{diam}(\mathbb{X})}{\epsilon\wedge 1}\right)+{d_{\mathbb{A}}}\log\left(1+\frac{3L\mathrm{diam}(\mathbb{A})}{\epsilon\wedge 1}\right)$
$\leq d_{\mathbb{X}}\log\left(1+3L\mathrm{diam}(\mathbb{X})\right)+{d_{\mathbb{A}}}\log\left(1+3L\mathrm{diam}(\mathbb{A})\right)+\frac{1}{2}(d_{\mathbb{X}}+d_{\mathbb{A}})\log n.$
Therefore, defining $D=d_{\mathbb{X}}\log\left(1+3L\mathrm{diam}(\mathbb{X})\right)+{d_{\mathbb{A}}}\log\left(1+3L\mathrm{diam}(\mathbb{A})\right)$, we conclude that by Proposition 4.1, the estimation error
@@
[…]